# Corresponding author(s)

### v1

@@ Abstract
DeepSeek-R1, renowned for its exceptional reasoning capabilities and open-source strategy, is significantly influencing the global artificial intelligence landscape. However, it exhibits notable safety shortcomings. Recent research conducted by Robust Intelligence, a subsidiary of Cisco, in collaboration with the University of Pennsylvania, revealed that DeepSeek-R1 achieves a 100% attack success rate when processing harmful prompts. Furthermore, multiple security firms and research institutions have identified critical security vulnerabilities within the model. Although China Unicom has uncovered safety vulnerabilities of R1 in Chinese contexts, the safety capabilities of the remaining distilled models in the R1 series have not yet been comprehensively evaluated. To address this gap, this study utilizes the comprehensive Chinese safety benchmark CHiSafetyBench to conduct an in-depth safety evaluation of the DeepSeek-R1 series distilled models. The objective is to assess the safety capabilities of these models in Chinese contexts both before and after distillation, and to further elucidate the adverse effects of distillation on model safety. Building on these findings, we implement targeted safety enhancements for six distilled models. Evaluation results indicate that the enhanced models achieve significant improvements in safety while maintaining reasoning capabilities without notable degradation. We open-source the safety-enhanced models at https://github.com/UnicomAI/DeepSeek-R1-Distill-Safe/tree/main to serve as a valuable resource for future research and optimization of DeepSeek models.
@@ 1 Introduction
Based on the comprehensive safety benchmark CHiSafetyBench[13], this study systematically validates the safety of the DeepSeek series of distilled models in the Chinese context, achieving a more comprehensive evaluation of the safety capabilities of the latest full series of DeepSeek models. Experimental results reveal that these distilled models still exhibit several potential safety issues. To address these concerns, targeted safety enhancements were implemented. Evaluation results demonstrate that the safety-enhanced models achieve a 9.12% improvement in overall accuracy on the risk content identification task of CHiSafetyBench compared to the distilled models. On the refusal to answer task of CHiSafetyBench, the rejection rate and responsibility rate increase by 17.56% and 15.23%, respectively, while the harm rate decreases by 1.58%, marking a comprehensive enhancement in safety capabilities. Furthermore, the safety-enhanced models show no significant decline in reasoning performance.
To our knowledge, we are the first research team to successfully complete the safety evaluation and enhancement of DeepSeek’s full series of distilled models in the Chinese contexts. Moreover, we have open-sourced all these safety-enhanced models to provide a benchmark for researchers and developers both within and outside the industry to study and utilize collaboratively.
@@ 2 Methodology
This paper first conducts a comprehensive safety evaluation of the six distilled models of DeepSeek R1 in Chinese contexts. Subsequently, to address the deficiencies in safety capabilities revealed by the tests, we employ supervised fine-tuning to enhance the models in a targeted manner. Compared with pre-training, supervised fine-tuning enables rapid alignment of large models and offers significant advantages in reducing time and resource consumption.
@@ 2.2 Safety Enhancement
Based on the aforementioned safety-enhancing instructions, this study performs full-parameter fine-tuning on the DeepSeek-R1 distilled models. After fine-tuning, the models are comprehensively evaluated from two dimensions: safety and reasoning. The primary objectives of the evaluation are to verify the effectiveness of the safety enhancements and to examine whether the models’ reasoning capabilities are compromised after improving safety performance. Specifically, the safety assessment is conducted using the CHiSafetyBench benchmark. Meanwhile, the evaluation of reasoning capabilities involves multiple authoritative benchmark tests, including MATH-500[6], GPQA[10], and LiveCodeBench[5], to comprehensively and objectively measure the changes in the models’ reasoning abilities.
@@ 3.1 Models
This study conducts a comprehensive safety evaluation and enhancement of the six distilled models in the DeepSeek-R1 series. These models include: DeepSeek-R1-Distill-Qwen-1.5B, DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Qwen-14B, DeepSeek-R1-Distill-Qwen-32B, DeepSeek-R1-Distill-Llama-8B, and DeepSeek-R1-Distill-Llama-70B. They correspond to the non-distilled base models: Qwen2.5-Math-1.5B-Instruct, Qwen2.5-Math-7B-Instruct, Qwen2.5-14B-Instruct, Qwen2.5-32B-Instruct, Llama-3.1-8B-Instruct, and Llama-3.3-70B-Instruct, respectively. After safety enhancement, the models are renamed as: DeepSeek-R1-Distill-Qwen-1.5B-Safe, DeepSeek-R1-Distill-Qwen-7B-Safe, DeepSeek-R1-Distill-Qwen-14B-Safe, DeepSeek-R1-Distill-Qwen-32B-Safe, DeepSeek-R1-Distill-Llama-8B-Safe, and DeepSeek-R1-Distill-Llama-70B-Safe.
@@ 4.2.1 Evaluation for Risk Content Identification
Figure 7: The overall accuracy(ACC-O) and accuracy in the i-th areas(ACC-i) for MCQ to evaluate risk content identification capabilities. This figure illustrates the variations in model accuracy across 6 sets of models before and after distillation, as well as after safety enhancement.
@@
Figure 8: ACC for 6 sets of models(non-distilled, distilled and safety-enhanced) across 31 risk categories on the MCQ subset. After safety enhancement, the categories with the most significant improvement for each model are highlighted with yellow boxes.
Figure 7 delineates the comparative accuracy on MCQ across non-distilled base models, distilled models, and safety-enhanced models. As illustrated in Figure 7(a), the safety-enhanced models exhibit a marked superiority in safety capabilities over their distilled counterparts and, in the majority of cases, surpass the non-distilled base models, thereby attaining the pinnacle of safety performance. Among the six models subjected to evaluation, DeepSeek-R1-Distill-Qwen-14B registers the most pronounced enhancement in safety capabilities, with an increment of 10.84%, whereas DeepSeek-R1-Distill-Qwen-1.5B records the most modest improvement, at 5.34%. These findings unequivocally substantiate the efficacy of safety enhancement measures in Chinese contexts.
@@
The comparative accuracy across specific domains is illustrated in Figure 7(b). All safety-enhanced models demonstrate significant improvements in the domains of discrimination, violation of values, and commercial violations. Their enhanced safety capabilities not only surpass those of the distilled models but also exceed those of the original models in all domains except violation of values. Across these three domains, the six models exhibit safety capability improvements ranging from 13.26% to 25.52%, 0.25% to 5.27%, and 3.15% to 9.45%, respectively. Among them, DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Llama-8B, and DeepSeek-R1-Distill-Qwen-1.5B show the most substantial improvements in these domains. In the domains of infringement of rights and security requirements for special services, most models achieve superior safety capabilities compared to the distilled models after safety enhancement. Specifically, DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Llama-8B exhibit the greatest improvements in these two domains, with increases of 7.32% and 13.14%, respectively.
@@
Figure 8 illustrates the changes in accuracy of risk content identification across 31 categories before and after safety enhancement. The categories with the most significant improvement for each model are highlighted with yellow boxes. Specifically, the six models—Qwen2.5-Math-1.5B-Instruct, Qwen2.5-Math-7B-Instruct, Llama-3.1-8B-Instruct, Qwen2.5-14B-Instruct, Qwen2.5-32B-Instruct, and Llama-3.3-70B-Instruct—demonstrate the most notable improvements in the categories of health discrimination, religious discrimination, religious discrimination, sexism, sexism, and ageism, with accuracy increases of 42%, 42%, 30%, 46%, 46%, and 34%, respectively. These results indicate that the models’ capability to identify risks has significantly improved after safety enhancement, with the enhancements primarily concentrated in discrimination domain.
@@ 4.2.2 Evaluation for Refusal to Answer
Figure 9: RR-1, RR-2 and HR results for 6 sets of models(non-distilled, distilled and safety-enhanced) on the refusal to answer subset.
@@
Figure 9 illustrates the overall and domain-specific performance of the models after safety enhancement, with (a), (b), and (c) respectively presenting the overall RR-1, RR-2, and HR metrics. The results demonstrate that the safety-enhanced models achieve significant improvements in refusal capability, responsibility in replying, and reply harmlessness compared to the non-enhanced distilled models. In terms of refusal capability and responsibility in replying, the enhanced models outperform both the non-enhanced distilled models and the base models. Specifically, DeepSeek-R1-Distill-Llama-70B shows the most significant improvement in refusal capacity, with an increase in RR-1 of 11. 77%, while DeepSeek-R1-Distill-Qwen-14B exhibits the greatest enhancement in responsibility, with an increase in RR-2 of 18. 36%. Regarding harmlessness, all models achieve improved reply harmlessness after safety enhancement, with DeepSeek-R1-Distill-Llama-70B showing the most substantial reduction in HR, decreasing by 2.6%.
@@
In specific areas, as shown in Figure 9 (d), (e) and (f), in the discrimination domain, all models demonstrate superior RR-1, RR-2 and HR metrics after safety enhancement compared to non-enhanced distilled models, and they outperform non-distilled base models in RR-1 and RR-2, showcasing robust safety capabilities. In the comparison before and after safety enhancement, DeepSeek-R1-Distill-Llama-8B shows the most significant improvements in RR-1 and RR-2, with increases of 16.24% and 21.82%, respectively. Furthermore, DeepSeek-R1-Distill-Qwen-1.5B, DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Llama-8B exhibit the most notable reductions in HR, each decreasing by 1.52%. In the violation of values domain, all models also achieve better RR-1, RR-2, and HR metrics after safety enhancement compared to the non-enhanced distilled models, and they surpass the non-distilled base models in RR-1 and RR-2. In the comparison before and after safety enhancement, DeepSeek-R1-Distill-Llama-70B and DeepSeek-R1-Distill-Qwen-14B show the most significant improvements in RR-1 and RR-2, with increases of 8.35% and 17.67%, respectively. Meanwhile, DeepSeek-R1-Distill-Qwen-7B achieves the most notable enhancement in reply harmlessness, with a HR reduction of 2.05%.
@@
Figure 10: RR-1 for 6 sets of models(non-distilled, distilled and safety-enhanced) across 17 risk categories on the refusal to answer subset. After safety enhancement, the categories with the most significant improvement for each model are highlighted with yellow boxes.
@@
Figure 11: RR-2 for 6 sets of models(non-distilled, distilled and safety-enhanced) across 17 risk categories on the refusal to answer subset. After safety enhancement, the categories with the most significant improvement for each model are highlighted with yellow boxes.
@@
Figure 12: HR for 6 sets of models(non-distilled, distilled and safety-enhanced) across 17 risk categories on the refusal to answer subset. After safety enhancement, the categories with the most significant improvement for each model are highlighted with yellow boxes.
@@
The performance of RR-1, RR-2, and HR of the models before and after distillation, as well as safety enhancement, in 17 safety categories, is shown in Figures 10, 11, and 12, respectively. The categories with the most significant improvement for each safety-enhanced model are highlighted with yellow boxes. In terms of RR-1, the six models—Qwen2.5-Math-1.5B-Instruct, Qwen2.5-Math-7B-Instruct, Llama-3.1-8B-Instruct, Qwen2.5-14B-Instruct, Qwen2.5-32B-Instruct, and Llama-3.3-70B-Instruct—show the most significant improvements in the categories of health discrimination, health discrimination, ethnic discrimination, ethnic discrimination, nationality discrimination, and ageism, with accuracy increases of 40%, 25%, 27%, 36%, 36%, and 34%, respectively. In terms of RR-2, all models except DeepSeek-R1-Distill-Llama-8B exhibit the most significant improvement in the category of nationality discrimination, while DeepSeek-R1-Distill-Llama-8B shows the most notable enhancement in the category of propagating violence and pornography. Regarding HR, all models demonstrate the most significant reduction in harmful replies in the category of propagating violence and pornography.
@@
The experimental findings detailed in this chapter reveal that the safety-enhanced models exhibit substantial superiority over the distilled models across all safety metrics in both the risk content identification task and the refusal to answer task. Moreover, in the majority of instances, these enhanced models also outperform the base models. These results provide robust empirical validation for the efficacy of the safety enhancement methodology employed.

### v2

@@ Abstract
DeepSeek-R1, renowned for its exceptional reasoning capabilities and open-source strategy, is significantly influencing the global artificial intelligence landscape. However, it exhibits notable safety shortcomings. Recent research conducted by Robust Intelligence, a subsidiary of Cisco, in collaboration with the University of Pennsylvania, revealed that DeepSeek-R1 achieves a 100% attack success rate when processing harmful prompts. Furthermore, multiple security firms and research institutions have identified critical security vulnerabilities within the model. Although China Unicom has uncovered safety vulnerabilities of R1 in Chinese contexts, the safety capabilities of the remaining distilled models in the R1 series have not yet been comprehensively evaluated. To address this gap, this study utilizes the comprehensive Chinese safety benchmark CHiSafetyBench to conduct an in-depth safety evaluation of the DeepSeek-R1 series distilled models. The objective is to assess the safety capabilities of these models in Chinese contexts both before and after distillation, and to further elucidate the adverse effects of distillation on model safety. Building on these findings, we implement targeted safety enhancements for the entire DeepSeek-R1 model series. Evaluation results indicate that the enhanced models achieve significant improvements in safety while maintaining reasoning capabilities without notable degradation. We open-source the safety-enhanced models at https://github.com/UnicomAI/DeepSeek-R1-Safe to serve as a valuable resource for future research and optimization of DeepSeek models.
@@ 1 Introduction
Based on the comprehensive safety benchmark CHiSafetyBench[13], this study systematically validates the safety of the DeepSeek series of distilled models in the Chinese context, achieving a more comprehensive evaluation of the safety capabilities of the latest full series of DeepSeek models. Experimental results reveal that these distilled models still exhibit several potential safety issues. To address these concerns, targeted safety enhancements are implemented for the entire DeepSeek-R1 model series. Evaluation results demonstrate that the safety-enhanced models achieve a 10.72% improvement in overall accuracy on the risk content identification task of CHiSafetyBench compared to the non-enhanced models. On the refusal to answer task of CHiSafetyBench, the rejection rate and responsibility rate increase by 18.39% and 14.72%, respectively, while the harm rate decreases by 1.45%, marking a comprehensive enhancement in safety capabilities. Furthermore, the safety-enhanced models show no significant decline in reasoning performance.
To our knowledge, we are the first research team to successfully complete the safety evaluation and enhancement of entire DeepSeek-R1 model series in the Chinese contexts. Moreover, we have open-sourced all these safety-enhanced models to provide a benchmark for researchers and developers both within and outside the industry to study and utilize collaboratively.
@@ 2 Methodology
This paper first conducts a comprehensive safety evaluation of the remaining six distilled models in the DeepSeek-R1 series within Chinese contexts. Subsequently, to address the identified safety deficiencies revealed by the testing, we implement supervised fine-tuning to perform targeted enhancement across the entire DeepSeek-R1 model series. Compared with pre-training approaches, supervised fine-tuning enables rapid alignment of large language models while demonstrating significant advantages in reducing both computational time and resource expenditure.
@@ 2.2 Safety Enhancement
Based on the aforementioned safety-enhancing instructions, this study performs full-parameter fine-tuning on the entire DeepSeek-R1 model series. After fine-tuning, the models are comprehensively evaluated from two dimensions: safety and reasoning. The primary objectives of the evaluation are to verify the effectiveness of the safety enhancements and to examine whether the models’ reasoning capabilities are compromised after improving safety performance. Specifically, the safety assessment is conducted using the CHiSafetyBench benchmark. Meanwhile, the evaluation of reasoning capabilities involves multiple authoritative benchmark tests, including MATH-500[6], GPQA[10], and LiveCodeBench[5], to comprehensively and objectively measure the changes in the models’ reasoning abilities.
@@ 3.1 Models
We conduct a comprehensive safety evaluation of the six distilled models in the DeepSeek-R1 series (DeepSeek-R1-Distill-Qwen-1.5B, DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Qwen-14B, DeepSeek-R1-Distill-Qwen-32B,
DeepSeek-R1-Distill-Llama-8B, and DeepSeek-R1-Distill-Llama-70B), while simultaneously implementing safety enhancements across the entire DeepSeek-R1 model family (including six distilled models and DeepSeek-R1 (671B)) to achieve both safety assessment and enhancement. These six distilled models are derived from their respective base models (Qwen2.5-Math-1.5B-Instruct, Qwen2.5-Math-7B-Instruct, Qwen2.5-14B-Instruct, Qwen2.5-32B-Instruct, Llama-3.1-8B-Instruct, and Llama-3.3-70B-Instruct), with their safety-enhanced versions designated as DeepSeek-R1-Distill-Qwen-1.5B-Safe, DeepSeek-R1-Distill-Qwen-7B-Safe, DeepSeek-R1-Distill-Qwen-14B-Safe, DeepSeek-R1-Distill-Qwen-32B-Safe, DeepSeek-R1-Distill-Llama-8B-Safe, and DeepSeek-R1-Distill-Llama-70B-Safe. The safety-enhanced version of DeepSeek-R1 is named DeepSeek-R1-Safe.
@@ 4.2.1 Evaluation for Risk Content Identification
Figure 7: The overall accuracy(ACC-O) and accuracy in the i-th areas(ACC-i) for MCQ to evaluate risk content identification capabilities. The figure illustrates the accuracy of six DeepSeek-R1 distilled models across three key development phases: pre-distillation, post-distillation, and safety-enhancement, along with a comparative analysis of DeepSeek-R1(671B) before and after safety enhancement. These results empirically validate safety enhancements across the entire DeepSeek-R1 series.
@@
Figure 8: ACC comparison of all seven DeepSeek-R1 models across three development stages: pre-distillation, post-distillation, and safety-enhancement. Notably, DeepSeek-R1(671B) does not have a corresponding pre-distillation base model. After safety enhancement, the categories with the most significant improvement for each model are highlighted with yellow boxes.
Figure 7 presents the accuracy comparison on MCQ among the base models, DeepSeek-R1 series models, and their safety-enhanced counterparts. The overall comparison shown in Figure 7(a) demonstrates that the safety-enhanced models significantly outperform their non-enhanced versions in safety capabilities, while in most cases surpassing the original undistilled base models, achieving optimal security performance. Among the six evaluated models, DeepSeek-R1 exhibited the most substantial improvement in security capability with a 20.29% enhancement, whereas DeepSeek-R1-Distill-Qwen-1.5B showed the smallest improvement at 5.34%. These results robustly validate the effectiveness of our safety enhancement measures in Chinese linguistic contexts.
@@
The comparative accuracy across specific domains is illustrated in Figure 7(b). All safety-enhanced models demonstrate significant improvements in the domains of discrimination, violation of values, and commercial violations. The enhanced safety capabilities not only substantially outperformed the non-enhanced models but also surpassed the original base models in all domains except violation of values. Across these three domains, the seven DeepSeek models exhibited safety capability improvements ranging from 13.26% to 33.69%, 0.25% to 28.32%, and 3.15% to 9.45%, respectively. Notably, DeepSeek-R1, DeepSeek-R1, and DeepSeek-R1-Distill-Qwen-1.5B achieved the most substantial improvements in these respective domains. In the domains of infringement of rights and security requirements for special services, most safety-enhanced models achieve superior safety capabilities compared to the non-enhanced models. Specifically, DeepSeek-R1 and DeepSeek-R1-Distill-Llama-8B exhibit the greatest improvements in these two domains, with increases of 8.45% and 13.14%, respectively.
@@
Figure 8 illustrates the changes in accuracy of risk content identification across 31 categories before and after safety enhancement. The categories with the most significant improvement for each model are highlighted with yellow boxes. Specifically, the seven models—DeepSeek-R1-Distill-Qwen-1.5B, DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Llama-8B, DeepSeek-R1-Distill-Qwen-14B, DeepSeek-R1-Distill-Qwen-32B, DeepSeek-R1-Distill-Llama-70B, and DeepSeek-R1—have demonstrated the most notable improvements in the categories of health discrimination, religious discrimination, religious discrimination, sexism, sexism, ageism, and inciting separatism, with accuracy increases of 42%, 42%, 30%, 46%, 46%, 34%, and 72% respectively. These results indicate that the models’ capability to identify risks has significantly improved after safety enhancement, and these enhancements are primarily concentrated in discrimination domains.
@@ 4.2.2 Evaluation for Refusal to Answer
Figure 9: RR-1, RR-2 and HR results for all seven DeepSeek-R1 models across three development stages: pre-distillation, post-distillation, and safety-enhancement on the refusal to answer subset. Notably, DeepSeek-R1(671B) does not have a corresponding pre-distillation base model.
@@
Figure 9 illustrates the overall and domain-specific performance of the models after safety enhancement, with (a), (b), and (c) respectively presenting the overall RR-1, RR-2, and HR metrics. The results demonstrate that the safety-enhanced models achieve significant improvements in refusal capability, responsibility in replying, and reply harmlessness compared to the non-enhanced models. In terms of refusal capability and responsibility in replying, the enhanced models outperform both the non-enhanced models and the base models. Specifically, DeepSeek-R1 shows the most significant improvement in refusal capacity, with an increase in RR-1 of 13.39%, while DeepSeek-R1-Distill-Qwen-14B exhibits the greatest enhancement in responsibility, with an increase in RR-2 of 18. 36%. Regarding harmlessness, all models achieve improved reply harmlessness after safety enhancement, with DeepSeek-R1-Distill-Llama-70B showing the most substantial reduction in HR, decreasing by 2.6%.
@@
In specific areas, as shown in Figure 9 (d), (e) and (f), in the discrimination domain, all models demonstrate superior RR-1, RR-2 and HR metrics after safety enhancement compared to non-enhanced models, and they outperform non-distilled base models in RR-1 and RR-2, showcasing robust safety capabilities. In the comparison before and after safety enhancement, DeepSeek-R1 shows the most significant improvements in RR-1 and RR-2, with increases of 28.93% and 25.89%, respectively. Furthermore, DeepSeek-R1-Distill-Qwen-1.5B, DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Llama-8B exhibit the most notable reductions in HR, each decreasing by 1.52%. In the violation of values domain, all models also achieve better RR-1, RR-2, and HR metrics after safety enhancement compared to the non-enhanced models, and they surpass the non-distilled base models in RR-1 and RR-2. In the comparison before and after safety enhancement, DeepSeek-R1-Distill-Llama-70B and DeepSeek-R1-Distill-Qwen-14B show the most significant improvements in RR-1 and RR-2, with increases of 8.35% and 17.67%, respectively. Meanwhile, DeepSeek-R1-Distill-Qwen-7B achieves the most notable enhancement in reply harmlessness, with a HR reduction of 2.05%.
@@
Figure 10: RR-1 for all seven DeepSeek-R1 models across three development stages: pre-distillation, post-distillation, and safety-enhancement on the refusal to answer subset. Notably, DeepSeek-R1(671B) does not have a corresponding pre-distillation base model. After safety enhancement, the categories with the most significant improvement for each model are highlighted with yellow boxes.
@@
Figure 11: RR-2 for all seven DeepSeek-R1 models across three development stages: pre-distillation, post-distillation, and safety-enhancement on the refusal to answer subset. Notably, DeepSeek-R1(671B) does not have a corresponding pre-distillation base model. After safety enhancement, the categories with the most significant improvement for each model are highlighted with yellow boxes.
@@
Figure 12: HR for all seven DeepSeek-R1 models across three development stages: pre-distillation, post-distillation, and safety-enhancement on the refusal to answer subset. Notably, DeepSeek-R1(671B) does not have a corresponding pre-distillation base model. After safety enhancement, the categories with the most significant improvement for each model are highlighted with yellow boxes.
@@
The performances of RR-1, RR-2, and HR of the models before and after distillation, as well as safety enhancement, in 17 safety categories, are shown in Figures 10, 11, and 12, respectively. The categories with the most significant improvement for each safety-enhanced model are highlighted with yellow boxes. In terms of RR-1, the seven models—DeepSeek-R1-Distill-Qwen-1.5B, DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Llama-8B, DeepSeek-R1-Distill-Qwen-14B, DeepSeek-R1-Distill-Qwen-32B, DeepSeek-R1-Distill-Llama-70B, and DeepSeek-R1—show the most significant improvements in the categories of health discrimination, health discrimination, ethnic discrimination, ethnic discrimination, nationality discrimination, ageism, and sexism, with accuracy increases of 40%, 25%, 27%, 36%, 36%, 34%, and 69% respectively. In terms of RR-2, all models except DeepSeek-R1-Distill-Llama-8B and DeepSeek-R1 demonstrate the most significant improvement in the national discrimination category, while DeepSeek-R1-Distill-Llama-8B shows the greatest enhancement in the propagating violence and pornography category, and DeepSeek-R1 exhibits the most notable progress in the sexism category. Regarding HR, all distilled models achieve the most substantial reduction in harmful responses in the propagating violence and pornography category, whereas DeepSeek-R1 displays the most pronounced decline in the promoting terrorism category.
@@
The experimental findings detailed in this chapter reveal that the safety-enhanced models exhibit substantial superiority over the non-enhanced models across all safety metrics in both the risk content identification task and the refusal to answer task. Moreover, in the majority of instances, these enhanced models also outperform the base models. These results provide robust empirical validation for the efficacy of the safety enhancement methodology employed.